# thanks: Equal contributorsthanks: Equal contributors

### v2

@@
thanks: Equal contributorsthanks: Equal contributors
Programming bistability in geometrically perturbed mechanical metamaterials
Yingchao Peng Viterbi School of Engineering, University of Southern California, CA 90089    Imtiar Niloy Department of Civil Engineering, Stony Brook University, NY 11794    Megan Kam Department of Civil Engineering, Stony Brook University, NY 11794    Paolo Celli paolo.celli@stonybrook.edu Department of Civil Engineering, Stony Brook University, NY 11794    Paul Plucinsky plucinsk@usc.edu Viterbi School of Engineering, University of Southern California, CA 90089
(January 17, 2024)
@@
Our work builds on these ideas under the lense of bistability and with particular emphasis on practical (reduced-order) design tools that can guide experiments at the conceptual/prototyping phase of design. We start by showing how to hard-encode bistability in a large class of periodic metamaterials. We then introduce an optimization framework for bistable metamaterials that incorporates a reduced-order model for the elastic energy, allowing us to tune the designs to achieve target morphing and elastic properties. A suite of representative examples and corresponding experiments follows. The experiments validate the bistable behavior of the patterns and certain qualitative features of their elastic properties, but also highlight how our theory can guide but not fully replace high-fidelity models and prototyping. Finally, we showcase the versatility of our design approach by extending it to non-periodic systems with complex shape-change.
@@
We begin by developing a general recipe for bistable planar kirigami comprised of a repeating unit cell of four quad panels and four quad slits, obtained through geometric perturbations of the rotating squares. The recipe amounts to a compact design formula for bistability, which we explain using Fig. 2 as a guide. Fig. 2A shows a generic unit cell representing the first stable state of a quad kirigami design, with the panels and slits labeled by 2D vectors $\mathbf{s}_{i}$, $\mathbf{t}_{i}$, $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$, $i=1,\ldots,4$. Fig. 2B illustrates its periodicity using the 2D Bravais lattice vectors $\bm{\ell}_{1}^{R}$ and $\bm{\ell}_{2}^{R}$. Fig. 2C shows the unit cell of the second stable state. Each panel in the cell is rotated in the plane by a right-hand rotation $\mathbf{R}(\phi_{i})$ of angle $\phi_{i}$, as shown, taking the initial cell vectors to deformed ones by the transformations $\mathbf{s}_{i}\mapsto\mathbf{R}(\phi_{i})\mathbf{s}_{i}$, $\mathbf{t}_{i}\mapsto\mathbf{R}(\phi_{i})\mathbf{t}_{i}$, $\mathbf{u}_{i}\mapsto\mathbf{R}(\phi_{i})\mathbf{u}_{i}$ and $\mathbf{v}_{i}\mapsto\mathbf{R}(\phi_{i})\mathbf{v}_{i}$, $i=1,\ldots,4$. Finally, Fig. 2D shows 2D Bravais lattice vectors $\bm{\ell}_{1}^{D}$ and $\bm{\ell}_{2}^{D}$, quantifying the periodicity in the second stable state. All of these vectors and rotation angles are subject to a variety of constraints for a compatible design, including equality constraints that enforce periodicity and ensure the vectors form closed loops about each panel and slit, as well as inequality constraints that ensure the panels are convex quadrilaterals that do not overlap. We enumerate all the compatibility conditions in SI Appendix, Section S1.A and manipulate them into forms broadly useful for design in SI Appendix, Sections S1.B and C.
@@
where $\mathbf{t},\mathbf{u}$ and $\mathbf{v}$ stack the corresponding $\mathbf{t}_{i},\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ design vectors into 8 component arrays, and $\bm{\ell}$ does likewise for the four Bravais lattice vectors $\bm{\ell}_{1}^{R},\ldots,\bm{\ell}_{2}^{D}$. The $28\times 12$ matrix $\mathbf{D}(\bm{\phi})$, concretely linking these arrays, is a lengthy nonlinear expression of the rotation angles $\bm{\phi}=(\phi_{1},\ldots,\phi_{4})$. Its explicit formula is provided in Eq. [S21] of SI Appendix, Section S1.B.
Eq. [1] organizes a wealth of information for designing and tuning bistable kirigami structures. The right-side contains the designer inputs. It includes the Bravais lattice vectors, which are the natural descriptors for target maximum stretch, Poisson’s ratio, and shearing between the two stable states. It also contains eight additional degrees of freedom (DOFs) through the four rotation angles $\bm{\phi}$ and design vectors $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$. Each can be tuned to achieve, for instance, a desirable energy barrier between the two stable states. The remaining parameters describing the designs in Fig. 2 are all stacked on the left-side of Eq. [1] and thus are fully determined from these designer inputs.
@@
An appealing aspect to this characterization is that it marries naturally with standard optimization tools to furnish a versatile design framework for tuning bistability. Assume a designer has in mind two stable states, obtained by prescribing the Bravais lattice vectors in the reference $\overline{\bm{\ell}_{1}^{R}},\overline{\bm{\ell}_{2}^{R}}$ and deformed $\overline{\bm{\ell}_{1}^{D}},\overline{\bm{\ell}_{2}^{D}}$ configurations. Since there are eight additional parameters on the right-side of Eq. [1], the design can be optimized to achieve any general objective that can be written as a minimization problem:
$\min~{}$$\big{\{}f_{\text{obj}}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})~{}\big{|}~{}\mathbf{g}_{\text{ineq}}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})\geq 0\big{\}}.$ (2)
In this formulation, $\mathbf{g}_{\text{ineq}}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})$ lists all the inequality constraints that are necessary and sufficient for the pattern to have convex panels and slits in its reference and deformed stable states. These constraints, which are written out explicitly in SI Appendix, Section S1.C, are nonlinear in all their arguments. Thus, Eq. [2] describes a constrained nonlinear optimization for which Matlab’s fmincon toolbox provides several well-developed and efficient numerical tools to find local minimizers. In other words, this optimization framework is ”ready-made” for engineering design.
We are particularly interested in objective functions that can assess and optimize a variety of features of the stored elastic energy of the kirigami design. The challenge is that calculating an elastic energy based on high fidelity modeling, like FEM or even bar-hinge based modeling [52], is not efficient and thus creates a bottleneck in the optimization process. We instead develop an elastic model that can be implemented directly into Matlab and evaluated using its fast solvers.
Our approach, formulated in detail in SI Appendix, Section S2.A, is illustrated in Fig. 2E. We model the corner points of the kirigami pattern as linear springs of zero rest length and unit stiffness and subject the pattern to the following simplified family of ”elastic” distortion: First, we rigidly rotate and translate each panel in an individual unit cell. Then, we repeat these panel motions cell by cell along the Bravais lattice vectors $\bm{\ell}_{1},\bm{\ell}_{2}$ indicated to achieve a bulk deformation of the pattern. This deformation elongates the springs, generating an elastic energy expressed in terms of the panel rotations and translations, and the Bravais lattice vectors.
After minimizing out the translations and lattice vectors in SI Appendix, Section S2.B, we obtain the revealing form for the energy
@@
for an $8\times 8$ symmetric and positive definite matrix $\mathbf{G}$ and the four panel rotations $\mathbf{R}(\eta_{1}),\ldots,\mathbf{R}(\eta_{4})$, as shown in the figure. The heuristics behind this energy are as follows. Prior to deformation, the pattern’s four types of slits satisfy $\sum_{i=1,\ldots,4}\mathbf{s}_{i}=\mathbf{0},\ldots,\sum_{i=1,\ldots,4}\mathbf{v}_{i}=\mathbf{0}$ because slits form closed loops. However, the deformed loops $\sum_{i=1,\ldots,4}\mathbf{R}(\eta_{i})\mathbf{s}_{i},\ldots,\sum_{i=1,\ldots,4}\mathbf{R}(\eta_{i})\mathbf{v}_{i}$ are typically broken ($\neq\mathbf{0})$ under the panel motions. Eq. [3] employs these broken loops as the fundamental measures of elastic strain in the pattern. The matrix $\mathbf{G}$ in this formula quantifies how the slits influence each other elastically. Its components range between values 0 and 1 independent of the krigami design, and are reported in Eq. [S36] of SI Appendix, Section S2.B.
@@
We call this energy the actuation energy. It depends kinematically only on the angle $\xi$ shown in Fig. 2F, describing the relative rotation between the first and second panel of each unit cell as the overall pattern is actuated. It is non-negative and satisfies $E_{\text{act}}=0$ when $\xi=0$ and $\xi=\phi_{2}-\phi_{1}$, reflecting the bistability hard encoded by Eq. [1].
@@
The first term tunes the designs so that the energy barrier between the designed stables states, $E_{b}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})=\max_{\xi\in(0,\phi_{2}-\phi_{1})}E_{\text{act}}(\xi,\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})$, is driven towards a specified target $E_{b}^{\text{targ}}\geq 0$, reflecting the amount of work needed to actuate the pattern from one stable state to the other. The second term optimizes for the stiffness $k_{1}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})=\partial_{\xi}\partial_{\xi}E_{\text{act}}(0,\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})/(\lambda^{\prime}(0))^{2}$ of the first stable state with respect to a characteristic stretch $\lambda(\xi)=\tfrac{|\mathbf{u}_{1}-\mathbf{v}_{1}+\mathbf{R}(\xi)(\mathbf{u}_{2}-\mathbf{v}_{2})|}{|\mathbf{u}_{1}-\mathbf{v}_{1}+\mathbf{u}_{2}-\mathbf{v}_{2}|}$ that takes the value $\lambda(0)=1$ in the first stable state and $\lambda(\phi_{2}-\phi_{1})=|\bm{\ell}_{1}^{D}|/|\bm{\ell}_{1}^{R}|$ in the second one (see SI Appendix, Section S2.D for more details). The designs are tuned by driving this stiffness towards a specified target $k_{1}^{\text{targ}}\geq 0$, allowing us to control whether we want this stable state to be ”locked-in” or have some give at its typical performance loads. Design tradeoffs are expected. For example, it is not usually possible to achieve a design that has both a high stiffness but a low overall energy barrier between the states. The numerical parameters $c_{b},c_{1}\geq 0$ express the desired importance of each term during an optimization. Though not done here, terms like the maximum force and/or the stiffness of the second stable state can also be included in the objective function.
@@
We illustrate the optimization framework by tuning kirigami designs to achieve a variety of energy barriers under a prototypical square-to-rectangle transformation. In the optimization, the Bravais lattice vectors are set at $\overline{\bm{\ell}_{1}^{R}}=\mathbf{e}_{1},\overline{\bm{\ell}_{2}^{R}}=\mathbf{e}_{2}$ and $\overline{\bm{\ell}_{1}^{D}}=1.2\mathbf{e}_{1}$, $\overline{\bm{\ell}_{2}^{D}}=0.8\mathbf{e}_{2}$ to encode the effectively square and rectangular stable states, and the moduli in Eq. [5] are taken as $c_{b}=1$ and $c_{1,2}=0$ to focus on optimizing for a target energy barrier. Fig. 3A-C shows three optimized designs, obtained by prescribing the target energy barrier from left to right as $E_{b}^{\text{targ}}=0.001,0.0015,0.003$ and performing the minimization in Eq. [2] in each case. A plot of actuation energy $E_{\text{act}}(\xi)$ versus stretch $\lambda(\xi)$ in Fig. 3E shows that each design achieves its target energy barrier. A randomly generated monostable design in Fig. 3D is included in Fig. 3E as another point of comparison.
@@
We propose three different fabrication strategies that differ in the way the energy barriers manifest as deformation, allowing us to explore the interplay between the fine design details inherent to prototyping and the theoretical predictions of bistability. In all cases, the specimens are made of 2$\times$2 unit cells and are tested in tension via a universal testing system using custom fixtures. Fig. 3F-H show the raw force-displacement curves for tension tests of the fabricated samples, as well as their stored energy curves (obtained by integration of the force curves). The color scheme for the curves distinguishes the different designs, just as in Fig. 3E. Additional details on the fabrication and experimental procedures are given in SI Appendix, Sections S4, S5. We only report experimental results in the main text; SI Appendix, Section S7 reports finite-element results that supplement these findings.
@@
In the first incarnation of our designs, we 3D print thick monolithic specimens made of ”soft” TPU-95, which represents the most conventional way to fabricate these metamaterials [15, 40, 45, 53]. Notably, each sample exhibits negligible out-of-plane deformations due to its large thickness, but also has elastic hinges that offer some resistance to the relative rotations between panels. Fig. 3F-1 shows the specimen with the largest (0.003) theoretical energy barrier in its first stable state; Fig. 3F-2 shows a snapshot of its in-plane deformation process during a tension test, with the distortions mostly concentrated in the hinge regions; Fig. 3F-3 shows the second stable state. This case is clearly bistable — the force-displacement curve dips below zero, resulting in two clear energy minima. However, bistability is far from guaranteed because hinge elasticity counteracts the geometric energy barriers that support a second stable state. In fact, the other two cases with the smaller theoretical energy barriers of 0.001 and 0.0015 are not bistable, as indicated by their force and energy curves.
@@
Our second fabrication strategy eliminates the hinge elasticity that opposes bistability, while keeping the actuation essentially planar. To do this, we 3D print ”skeletal” and bi-material panels and assemble them via actual pin joints. The printed panels are composed of soft TPU bars and stiff nylon hinge regions to ensure that the deformations concentrate in the bars rather than in the neighborhood of the pin joints. We also place the specimens between clear acrylic plates during testing to prevent out-of-plane deformation. Fig. 3G-1 shows the 0.003 design fabricated in this fashion in its first stable state; Fig. 3G-2 shows an intermediate state during testing, illustrating how the deformation within the panels manifests as in-plane bending and buckling of the bars; Fig. 3G-3 shows the second stable state. The experimental force-displacement curves for this specimen clearly indicate bistability. Note that the presence of pin-joints allows the specimen to be stress-free in its second stable state even though the curve shows that the energy is non-zero in this state. We attribute the tilted energy curve to friction and material viscoelasticity which dissipate energy during the tests. While this second incarnation is a better candidate to demonstrate bistability for a broad range of designs, fabrication proved challenging and time consuming; for this reason, we only report results for the 0.003 specimen.
@@
The final incarnation also features pin joints, but the panels are now much thinner and laser cut out of PETG. As shown in Fig. 3H-1 to H-3 for the 0.003 sample, these specimens transition between stable states via out-of-plane bending of the panels. The force and energy curves show that the patterns behave as expected — all the theoretically bistable designs are indeed bistable and the magnitude of their energy barriers trends with that of the theory. Once again, frictional losses cause the energy curves to display non-zero values at the second equilibrium even though these states are stress-free.
@@
Overall, these case studies validate our purely geometric design and optimization tools, as they showcase a variety of bistable metamaterials whose shape change matches the theory and whose energy barriers match the trends of the theory. We envision that the synergy between optimization and prototyping can be improved by introducing ”non-universal” features into the objective function that depend on the choice of fabrication strategy, although we do not pursue this further.
@@
Figure 4: Design explorations. (A) Matrix of designs optimized to attain axial shape transformation (from a square to another square or rectangle) of various degree, with a target energy barrier of 0.003. Designs are color-coded according to the actual value of the barrier achieved during optimization, as indicated in the legend below (B). (B) Suite of designs that undergo a shear-type shape transformation. $\lambda_{1}$, $\lambda_{2}$ and $\gamma$, which control the magnitude of the shape change, are defined in Eq. 6. (C-E) First and second stable states of various patterns, with their mono-material physical realizations. (C) Pattern designed to undergo extreme shear-type morphing, with $\gamma=0.8$; (D) Extreme non-auxetic morphing; (E) Extreme auxetic morphing. In all experimental images, scale bar: 5 cm. (F) Theoretical energy landscapes of patterns designed to display various target stiffness $k_{1}^{\text{targ}}$, while maintaining a barrier around 0.001. In the legend, we show the first stable states of all such patterns.

### v3

@@
Abstract
@@
This article may be downloaded for personal use only. Any other use requires prior permission of the author and the American Physical Society. This article appeared in: Physical Review Applied 22, 014073 (2024) and may be found at: https://doi.org/10.1103/PhysRevApplied.22.014073
@@
Our work builds on these ideas under the lense of bistability and with particular emphasis on practical (reduced-order) design tools that can guide experiments at the conceptual/prototyping phase of design. We start by showing how to hard-encode bistability in a large class of 2D periodic metamaterials composed of repeating unit cells of panels and slits, termed planar kirigami herein and elsewhere [45, 52, 53], as opposed to kirigami that rely on out-of-plane buckling [54, 55]. We then introduce an optimization framework for bistable planar kirigami that incorporates a reduced-order model for the elastic energy, allowing us to tune the designs to achieve target morphing and elastic properties. A suite of representative examples and corresponding experiments follows. We explore examples ranging from classical monolithic planar kirigami configurations to pin-jointed panel systems and truss-based analogs to test the applicability of our design and optimization strategies in a variety of settings. In each case, the experiments validate the bistable behavior of the patterns and certain qualitative features of their elastic properties, but also highlight how our theory can guide but not fully replace high-fidelity models and prototyping. Finally, we showcase the versatility of our design approach by extending it to non-periodic systems with complex shape-change.
@@
We begin by developing a general recipe for bistable planar kirigami comprised of a repeating unit cell of four quad panels and four quad slits, obtained through geometric perturbations of the rotating squares. The recipe amounts to a compact design formula for bistability, which we explain using Fig. 2 as a guide. Fig. 2A shows a generic unit cell representing the first stable state of a quad kirigami design, with the panels and slits labeled by 2D vectors $\mathbf{s}_{i}$, $\mathbf{t}_{i}$, $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$, $i=1,\ldots,4$. Fig. 2B illustrates its periodicity using the 2D Bravais lattice vectors $\bm{\ell}_{1}^{R}$ and $\bm{\ell}_{2}^{R}$. Fig. 2C shows the unit cell of the second stable state. Each panel in the cell is rotated in the plane by a right-hand rotation $\mathbf{R}(\phi_{i})$ of angle $\phi_{i}$, as shown, taking the initial cell vectors to deformed ones by the transformations $\mathbf{s}_{i}\mapsto\mathbf{R}(\phi_{i})\mathbf{s}_{i}$, $\mathbf{t}_{i}\mapsto\mathbf{R}(\phi_{i})\mathbf{t}_{i}$, $\mathbf{u}_{i}\mapsto\mathbf{R}(\phi_{i})\mathbf{u}_{i}$ and $\mathbf{v}_{i}\mapsto\mathbf{R}(\phi_{i})\mathbf{v}_{i}$, $i=1,\ldots,4$. Finally, Fig. 2D shows 2D Bravais lattice vectors $\bm{\ell}_{1}^{D}$ and $\bm{\ell}_{2}^{D}$, quantifying the periodicity in the second stable state. All of these vectors and rotation angles are subject to a variety of constraints for a compatible design, including equality constraints that enforce periodicity and ensure the vectors form closed loops about each panel and slit, as well as inequality constraints that ensure the panels are convex quadrilaterals that do not overlap. We enumerate all the compatibility conditions in Supplemental Material, Section S1.A and manipulate them into forms broadly useful for design in Supplemental Material, Sections S1.B and C [56].
@@
where $\mathbf{t},\mathbf{u}$ and $\mathbf{v}$ stack the corresponding $\mathbf{t}_{i},\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ design vectors into 8 component arrays, and $\bm{\ell}$ does likewise for the four Bravais lattice vectors $\bm{\ell}_{1}^{R},\ldots,\bm{\ell}_{2}^{D}$. The $28\times 12$ matrix $\mathbf{D}(\bm{\phi})$, concretely linking these arrays, is a lengthy nonlinear expression of the rotation angles $\bm{\phi}=(\phi_{1},\ldots,\phi_{4})$. Its explicit formula is provided in Eq. [S21] of Supplemental Material, Section S1.B [56].
Eq. [LABEL:eq:theorem1] organizes a wealth of information for designing and tuning bistable kirigami structures. The right-side contains the designer inputs. It includes the Bravais lattice vectors, which are the natural descriptors for target maximum stretch, Poisson’s ratio, and shearing between the two stable states. It also contains eight additional degrees of freedom (DOFs) through the four rotation angles $\bm{\phi}$ and design vectors $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$. Each can be tuned to achieve, for instance, a desirable energy barrier between the two stable states. The remaining parameters describing the designs in Fig. 2 are all stacked on the left-side of Eq. [LABEL:eq:theorem1] and thus are fully determined from these designer inputs.
@@
An appealing aspect to this characterization is that it marries naturally with standard optimization tools to furnish a versatile design framework for tuning bistability. Assume a designer has in mind two stable states, obtained by prescribing the Bravais lattice vectors in the reference $\overline{\bm{\ell}_{1}^{R}},\overline{\bm{\ell}_{2}^{R}}$ and deformed $\overline{\bm{\ell}_{1}^{D}},\overline{\bm{\ell}_{2}^{D}}$ configurations. Since there are eight additional parameters on the right-side of Eq. [LABEL:eq:theorem1], the design can be optimized to achieve any general objective that can be written as a minimization problem:
$\min{}$$\big{\{}f_{\text{obj}}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})~{}\big{|}~{}\mathbf{g}_{\text{ineq}}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})\geq 0\big{\}}.$ (2)
In this formulation, $\mathbf{g}_{\text{ineq}}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})$ lists all the inequality constraints that are necessary and sufficient for the pattern to have convex panels and slits in its reference and deformed stable states. These constraints, which are written out explicitly in SI Appendix, Section S1.C, are nonlinear in all their arguments. Thus, Eq. [2] describes a constrained nonlinear optimization for which Matlab’s fmincon toolbox provides several well-developed and efficient numerical tools to find local minimizers. In other words, this optimization framework is ”ready-made” for engineering design. See the flowchart in Fig.S3 and Supplemental Material, Section S2.E [56] for additional details on numerical aspects of this framework.
All that remains now is to prescribe an objective function for the optimization. We are particularly interested in objective functions that can assess and optimize a variety of features of the stored elastic energy of the kirigami design. The challenge is that calculating an elastic energy based on high fidelity modeling, like FEM or even bar-hinge based modeling [57], is not efficient and thus creates a bottleneck in the optimization process. We instead develop an elastic model that can be implemented directly into Matlab and evaluated using its fast solvers.
Our approach is formulated in detail in Supplemental Material, Section S2.A [56] and illustrated in Fig. 2E. We model the corner points of the kirigami pattern as linear springs of zero rest length and unit stiffness and allow the panels to rotate and translate by a periodic motion that matches a bulk deformation expressed by the Bravais lattice vectors $\bm{\ell}_{1},\bm{\ell}_{2}$ in the figure. This deformation elongates the springs, generating an elastic energy expressed in terms of the panel rotations and translations, and the Bravais lattice vectors.
After minimizing out the translations and lattice vectors in Supplemental Material, Section S2.B [56], we obtain the revealing form for the energy
@@
for an $8\times 8$ symmetric and positive definite matrix $\mathbf{G}$ and the four panel rotations $\mathbf{R}(\eta_{1}),\ldots,\mathbf{R}(\eta_{4})$, as shown in the figure. The heuristics behind this energy are as follows. Prior to deformation, the pattern’s four types of slits satisfy $\sum_{i=1,\ldots,4}\mathbf{s}_{i}=\mathbf{0},\ldots,\sum_{i=1,\ldots,4}\mathbf{v}_{i}=\mathbf{0}$ because slits form closed loops. However, the deformed loops $\sum_{i=1,\ldots,4}\mathbf{R}(\eta_{i})\mathbf{s}_{i},\ldots,\sum_{i=1,\ldots,4}\mathbf{R}(\eta_{i})\mathbf{v}_{i}$ are typically broken ($\neq\mathbf{0})$ under the panel motions. Eq. [LABEL:eq:springEnergyMain] employs these broken loops as the fundamental measures of elastic strain in the pattern. The matrix $\mathbf{G}$ in this formula quantifies how the slits influence each other elastically. Its components range between values 0 and 1 independent of the krigami design, and are reported in Eq. [S36] of Supplemental Material, Section S2.B [56].
@@
We call this energy the actuation energy. It depends kinematically only on the angle $\xi$ shown in Fig. 2F, describing the relative rotation between the first and second panel of each unit cell as the overall pattern is actuated. It is non-negative and satisfies $E_{\text{act}}=0$ when $\xi=0$ and $\xi=\phi_{2}-\phi_{1}$, reflecting the bistability hard encoded by Eq. [LABEL:eq:theorem1].
@@
The first term tunes the designs so that the energy barrier between the designed stables states, $E_{b}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})=\max_{\xi\in(0,\phi_{2}-\phi_{1})}E_{\text{act}}(\xi,\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})$, is driven towards a specified target $E_{b}^{\text{targ}}\geq 0$, reflecting the amount of work needed to actuate the pattern from one stable state to the other. The second term optimizes for the stiffness $k_{1}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})=\partial_{\xi}\partial_{\xi}E_{\text{act}}(0,\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})/(\lambda^{\prime}(0))^{2}$ of the first stable state with respect to a characteristic stretch $\lambda(\xi)=\tfrac{|\mathbf{u}_{1}-\mathbf{v}_{1}+\mathbf{R}(\xi)(\mathbf{u}_{2}-\mathbf{v}_{2})|}{|\mathbf{u}_{1}-\mathbf{v}_{1}+\mathbf{u}_{2}-\mathbf{v}_{2}|}$ that takes the value $\lambda(0)=1$ in the first stable state and $\lambda(\phi_{2}-\phi_{1})=|\bm{\ell}_{1}^{D}|/|\bm{\ell}_{1}^{R}|$ in the second one (see Supplemental Material, Section S2.D [56] for more details). The designs are tuned by driving this stiffness towards a specified target $k_{1}^{\text{targ}}\geq 0$, allowing us to control whether we want this stable state to be ”locked-in” or have some give at its typical performance loads. Design tradeoffs are expected. For example, it is not usually possible to achieve a design that has both a high stiffness but a low overall energy barrier between the states. The numerical parameters $c_{b},c_{1}\geq 0$ express the desired importance of each term during an optimization. Though not done here, terms like the maximum force and/or the stiffness of the second stable state can also be included in the objective function.
@@
We illustrate the optimization framework by tuning kirigami designs to achieve a variety of energy barriers under a prototypical square-to-rectangle transformation. In the optimization, the Bravais lattice vectors are set at $\overline{\bm{\ell}_{1}^{R}}=\mathbf{e}_{1},\overline{\bm{\ell}_{2}^{R}}=\mathbf{e}_{2}$ and $\overline{\bm{\ell}_{1}^{D}}=1.2\mathbf{e}_{1}$, $\overline{\bm{\ell}_{2}^{D}}=0.8\mathbf{e}_{2}$ to encode the effectively square and rectangular stable states, and the moduli in Eq. [5] are taken as $c_{b}=1$ and $c_{1}=0$ to focus on optimizing for a target energy barrier. Fig. 3A-C shows three optimized designs, obtained by prescribing the target energy barrier from left to right as $E_{b}^{\text{targ}}=0.001,0.0015,0.003$ and performing the minimization in Eq. [2] in each case. A plot of actuation energy $E_{\text{act}}(\xi)$ versus stretch $\lambda(\xi)$ in Fig. 3E shows that each design achieves its target energy barrier. A randomly generated monostable design in Fig. 3D is included in Fig. 3E as another point of comparison.
@@
We propose three different fabrication strategies that differ in the way the energy barriers manifest as deformation, allowing us to explore the interplay between the fine design details inherent to prototyping and the theoretical predictions of bistability. In all cases, the specimens are made of 2$\times$2 unit cells and are tested in tension via a universal testing system using custom fixtures. Fig. 3F-H show the raw force-displacement curves for tension tests of the fabricated samples, as well as their stored energy curves (obtained by integration of the force curves). The color scheme for the curves distinguishes the different designs, just as in Fig. 3E. Additional details on the fabrication and experimental procedures, and on the dimensions of specimens, are given in Supplemental Material, Sections S4, S5 [56]. We only report experimental results in the main text; Supplemental Material, Section S7 [56] reports finite-element results that complement these findings.
@@
In the first incarnation of our designs, we 3D print thick monolithic specimens made of ”soft” TPU-95, which represents the most conventional way to fabricate these metamaterials [15, 40, 45, 58]. Notably, each sample exhibits negligible out-of-plane deformations due to its large thickness, but also has elastic hinges that offer some resistance to the relative rotations between panels. Fig. 3F-1 shows the specimen with the largest (0.003) theoretical energy barrier in its first stable state; Fig. 3F-2 shows a snapshot of its in-plane deformation process during a tension test, with the distortions mostly concentrated in the hinge regions; Fig. 3F-3 shows the second stable state, which is qualitatively similar to the theoretical one, while featuring some localized bending near the hinges. This case is clearly bistable — the force-displacement curve dips below zero, resulting in two clear energy minima. However, bistability is far from guaranteed because hinge elasticity counteracts the geometric energy barriers that support a second stable state. In fact, the other two cases with the smaller theoretical energy barriers of 0.001 and 0.0015 are not bistable, as indicated by their force and energy curves.
@@
Our second fabrication strategy eliminates the hinge elasticity that opposes bistability, while keeping the actuation essentially planar. To do this, we 3D print “skeletal” and bi-material panels and assemble them via actual pin joints. The printed panels are composed of soft TPU bars and stiff nylon hinge regions to ensure that the deformations concentrate in the bars rather than in the neighborhood of the pin joints. We also place the specimens between clear acrylic plates during testing to prevent out-of-plane deformation. Fig. 3G-1 shows the 0.003 design fabricated in this fashion in its first stable state; Fig. 3G-2 shows an intermediate state during testing, illustrating how the deformation within the panels manifests as in-plane bending and buckling of the bars; Fig. 3G-3 shows the second stable state, which now matches the theoretical one. The experimental force-displacement curves for this specimen clearly indicate bistability. Note that the presence of pin-joints allows the specimen to be stress-free in its second stable state even though the curve shows that the energy is non-zero in this state. We attribute the tilted energy curve to friction and material viscoelasticity, which dissipate energy during the tests. While this second incarnation is a better candidate to demonstrate bistability for a broad range of designs, fabrication proved challenging and time consuming; for this reason, we only report results for the 0.003 specimen.
@@
The final incarnation also features pin joints, but the panels are now much thinner and laser cut out of PETG. As shown in Fig. 3H-1 to H-3 for the 0.003 sample, these specimens transition between stable states via out-of-plane bending of the panels. The force and energy curves show that the patterns behave as expected — all the theoretically bistable designs are indeed bistable and the magnitude of their energy barriers trends with that of the theory. In particular, the energy barrier for the 0.003 specimen is larger than the 0.0015 one, which is in turn larger than the 0.001 case. Here, significant frictional losses due to the panels pushing against each other and the rivets during out-of-plane deformation cause the energy curves to display non-zero values at the second equilibrium, even though these states are stress-free.
@@
Overall, these case studies validate our purely geometric design and optimization tools, as they showcase a variety of bistable metamaterials whose shape change matches the theory and whose energy barriers match the trends of the theory. We envision that the synergy between optimization and prototyping can be improved by introducing “non-universal” features into the objective function (Eq. [5]) that depend on the choice of fabrication strategy, although we do not pursue this further.
@@
Figure 4: Design explorations. (A) Matrix of designs optimized to attain axial shape transformation (from a square to another square or rectangle) of various degree, with a target energy barrier of 0.003. Designs are color-coded according to the actual value of the barrier achieved during optimization, as indicated in the legend below (B). Shaded designs in the lower diagonal are simply rotations of the corresponding designs in the upper diagonal. (B) Suite of designs that undergo a shear-type shape transformation. $\lambda_{1}$, $\lambda_{2}$ and $\gamma$, which control the magnitude of the shape change, are defined in Eq. 6. (C-E) First and second stable states of various patterns, with their mono-material physical realizations. (C) Pattern designed to undergo extreme shear-type morphing, with $\gamma=0.8$; (D) Extreme non-auxetic morphing; (E) Extreme auxetic morphing. In all experimental images, scale bar: 5 cm.